OpenStudy (anonymous):

Question #3: Write the inequality in SLOPE INTERCEPT FORM 9x-3y≤-12 <3

OpenStudy (johnweldon1993):

ahhh this ones different here.... same steps 9x - 3y <= -12 -3y <= -12 - 9x ********you want to divide by -3 to isolate y right? when dividing or multiplying by a negative number in inequalities....SWITCH THE DIRECTION OF THE SIGN!***** so -3y <= -12 - 9x becomes y >=!!!!! 4 + 9/3x written correctly \[y \ge 9/3x + 4\] :)